OpenStudy (anonymous):

change y^2-4x^2-16x-2y-19=0 to standard form

OpenStudy (anonymous):

\[\begin{align*} y^2-2y-4x^2-16x-19&=0\\ \left(y^2-2y+1\right)-4\left(x^2+4x\right)-20&=0\\ \left(y-1\right)^2-4\left(x^2+4x+4-4\right)-20&=0\\ \left(y-1\right)^2-4\left((x+2)^2-4\right)-20&=0\\ \left(y-1\right)^2-4(x+2)^2+16-20&=0\\ \left(y-1\right)^2-4(x+2)^2&=4\\ \frac{\left(y-1\right)^2}{4}-(x+2)^2&=1\\ \end{align*}\]